OpenStudy (braves12):

An Indian tepee is a conical tent. Find the number of cubic feet of air in a teepee 10 ft. in diameter and 12 ft. high. 100 cu. ft. 300 cu. ft. 400 cu. ft.

OpenStudy (anonymous):

the volume of anything pointy at the top and with straight sides is (1/3) height * base area here height = 12 base area = pi (5)^2 = 25 pi or about 75 so (1/3)*12 * 75 = about 300